5. (4 points) When a particular reaction is run at a constant pressure, 1.7 kJ of work is done by the reaction and 30.6 k.J of energy is absorbed by the reaction as heat. Calculate ??,xn .
Answer #1

work is done by the reaction and thus it is positive i.e, W = + 1.7* 103 J

Heat energy is absorbed by the reaction thus it is positive i.e, Q = 30.6 * 103 J

Change in internal energy = Q + W, = 1.7* 103 J + 30.6 * 103 J ; = 32.3 KJ